OpenStudy (anonymous):

Solve 4x - 2x - 2 = -42. A. -20 B. -7 C. -5 D. 0

OpenStudy (anonymous):

the answer is A

sam (.sam.):

4x-2x-2=-42 Since 4x and -2x are like terms, add -2x to 4x to get 2x. 2x-2=-42 Since -2 does not contain the variable to solve for, move it to the right-hand side of the equation by adding 2 to both sides. 2x=2-42 Subtract 42 from 2 to get -40. 2x=-40

OpenStudy (anonymous):

Thank you guy's very much.

OpenStudy (anonymous):

ur welcome:)

sam (.sam.):

2x=-40 divide 2 to both sides and get your answer

OpenStudy (anonymous):

x=-20 So A